# VisionReward: Fine-Grained Multi-Dimensional Human Preference Learning for Image and Video Generation

### v1

@@ Abstract
We present a general strategy to aligning visual generation models—both image and video generation—with human preference. To start with, we build VisionReward—a fine-grained and multi-dimensional reward model. We decompose human preferences in images and videos into multiple dimensions, each represented by a series of judgment questions, linearly weighted and summed to an interpretable and accurate score. To address the challenges of video quality assessment, we systematically analyze various dynamic features of videos, which helps VisionReward surpass VideoScore by 17.2% and achieve top performance for video preference prediction. Based on VisionReward, we develop a multi-objective preference learning algorithm that effectively addresses the issue of confounding factors within preference data. Our approach significantly outperforms existing image and video scoring methods on both machine metrics and human evaluation. All code and datasets are provided at https://github.com/THUDM/VisionReward.
@@ 1 Introduction
Figure 1: Samples of VisionReward and Multi-Objective Preference Optimization (MPO) algorithm.
Visual generation models, including text-to-image [8, 30, 32, 31, 2, 27] and text-to-video [15, 14, 39, 51, 4, 47] generation, have developed rapidly in recent times. Given a textual description, these visual generation models can produce high-quality images or videos. Inspired by reinforcement learning from human feedback (RLHF) in large language model (LLM) [26], researchers have found it crucial to align text-to-image with human preference [46], due to various challenges such as alignment, fidelity, aesthetic and safety. A series of works [46, 20, 44] have explored training reward models that simulate human preferences, and subsequent works [3, 9, 6, 45, 40] have developed optimization methods for diffusion models based on these reward models or preference dataset. With the development of text-to-video models, similar evaluation [17, 11] and optimization [48, 28] methods have been attempted and get improvement.
Despite advancements, current RLHF methods for text-to-vision models still face substantial challenges:
  • Reward Models are Biased and Inexplicable. Current reward models learn from human preference, which contain many trade-offs between different factors, leading to preference biases.  Fig. 1 (a) demonstrates an example.
  • Evaluation for Video is Challenging. It’s difficult to assess dynamic quality of videos, such as movement reality and motion smoothness, as shown in  Fig. 1 (b).
  • Over-optimization or Lack-optimization. Existing RLHF methods tend to over-optimize or weaken certain factors, resulting in suboptimal outcomes after optimization, as revealed in  Fig. 1 (c)(d).
Contributions. To address these challenges, we propose a fine-grained, multi-dimensional reward model for text-to-vision generation—VisionReward—which performs a variety of judgmental questions and answers, and uses linear weighting to predict human preference. Based on VisionReward, we develop MPO algorithm to stably optimize visual generation models. Our contributions include:
  • We design a unified annotation system for both image and video generation, decomposing the factors influencing human preferences. To address the challenges of video evaluation, we incorporate extensive observations of dynamic content in videos into our judgment tasks, such as motion stability or movement quality. The annotation contains 3 million questions for 48k images and 2 million questions for 33k videos. This dataset enables a unified training pipeline for VisionReward (shown as  Fig. 2).
  • We demonstrate that VisionReward achieves high accuracy and is interpretable in predicting human preferences. VisionReward outperforms existing methods for preference prediction, especially on video assessment, surpassing VideoScore [11] by 17.2 %.
  • We introduce Multi-Objective Preference Optimization (MPO) for stably tuning visual generation models, avoiding over-optimization or lack of optimization for certain factors. MPO with VisionReward outperforms directly tuned with human annotation or other RM.
2 Method
2.1 Annotation
@@
Fine-grained Questions Design. When evaluating an image or video, human preferences are often a result of the interplay of multiple factors, necessitating a balance among various considerations. To deconstruct human preferences systematically, we develop a fine-grained and multi-dimensional preference decomposition framework.  Tab. 1 illustrates our preference dimensions and the checklist count. We identify 5 dimensions for text-to-image generation and expand to 9 dimensions for text-to-video.
In our annotation process, we design multiple-choice questions for each dimension, allowing annotators to select the option that best fits the image. Additionally, we further decompose these options into several binary (yes/no) questions, ultimately forming a checklist for evaluating each image or video. This checklist is then utilized for subsequent evaluation systems and reward model training.
Annotation Data Preparation. To ensure diversity in our annotated data, we sample from various data sources, as presented in  Tab. 2. Following this fine-grained annotation process, we compile a dataset comprising 48k images and 33k videos.
Dimension #Sub-dimension #Checklist
Image Video Image Video
Alignment 1 1 1 4
Composition 5 1 13 2
Quality 5 4 14 14
Fidelity 5 3 25 9
Safety&Emotion 2 1 8 4
Stability - 5 - 12
Dynamic - 2 - 8
Physics - 1 - 4
Preservation - 2 - 7
Total 18 20 61 64
Table 1: Taxonomy of annotation for VisionReward.
Type Source #Samples #Checklist
Image ImageRewardDB [46] 16K 1M
Pick-a-Pic [20] 16K 1M
HPDv2 [44] 16K 1M
Video CogVideoX [47] 10K 0.6M
Open-Sora [51] 10K 0.6M
VideoCrafter2 [4] 10K 0.6M
Panda-70M [5] 3K 0.2M
Table 2: Statistics of source data and annotation.
Figure 2: An overview of the VisionReward and Multi-Objective Preference Optimization (MPO).
2.2 VisionReward Training
VisionReward consists of two components:
  1. 1.
    Checklist: VisionReward is a vision-language generative model to answer a set of judgment questions regarding an image or video with ”yes” or ”no”.
  2. 2.
    Linear Weighted Summarize: VisionReward applies a linear weighting to the list of ”yes” or ”no” responses to obtain the final score.
@@ 2.2 VisionReward Training
Checklist. Specifically, we use CogVLM2 [16] as the base model for image understanding, and CogVLM2-Video as the base model for video understanding. In terms of data, we have obtained millions of annotated binary judgment results. Initially, we performed a balanced sampling on each judgment question by addressing the imbalance between positive and negative examples, ensuring a roughly equal number of positive and negative instances associated with each judgment question. Then we use balanced instruction tuning dataset consisting of judgement questions to fine-tune base VLM.
@@
Regression. Based on a series of binary responses (yes or no), we use linear weighted summaries to predict human preferences. This is achieved by mapping yes and no to 1 and -1, respectively, and constructing a feature vector $\mathbf{X}=\{x^{i}\}$, where $i=1,\ldots,n$ and $n$ represents the number of binary questions. We use a set of linear weights to obtain a score that accurately reflects human preferences, expressed as follows:
@@
Through minimizing this objective function, we aim to find the optimal weights $\mathbf{w}$ that best represent human preferences.
2.3 Multi-Objective Preference Optimization
@@
Challenges. Although Diffusion-DPO has designed a direct preference optimization method to fine-tune text-to-image diffusion models, it suffers from being biased and over-optimized on certain dimensions such as aesthetic but gets somewhat worse, for example, less safe and more trouble about human body. We reproduce the procedure of Diffusion-DPO and test the fine-tuned model on 10,000 sampled prompts from training data, using VisionReward to score images.  Fig. 3 illustrates the changes across various dimensions after applying preference learning using the human preference dataset in which factors are coupled with different trade-offs. Formally, let the factors influencing human preferences be $D=\{d_{1},d_{2},\ldots,d_{n}\}$. During preference learning, it is likely that when there is a trade-off between $d_{i}$ and $d_{j}$, humans may prioritize $d_{i}$ more, leading to an increase in $d_{i}$ and a decrease in $d_{j}$.
@@
MPO: Insight and Solution. We thus propose an algorithm (Cf. Algorithm 1) to ensure that all dimensions are reasonably enhanced. We merge questions of VisionReward into these dimensions, formulated as reward $R=R(d_{k})$ where $k\in\{1,\ldots,n\}$ and $R(d_{k})=\sum(\text{reward}(q))$ with $q\in d_{k}$. Given two images $x^{i}$ and $x^{j}$, with their corresponding preference factor vectors being $\{d_{1}^{i},\ldots,d_{n}^{i}\}$ and $\{d_{1}^{j},\ldots,d_{n}^{j}\}$ respectively, we define a situation where $R^{i}$ dominates $R^{j}$ if $R(d_{k}^{i})\geq R(d_{k}^{j})$ for any $k\in\{1,\ldots,n\}$. We can use the dominant pairs to ensure that each dimension $d$ is not weakened in the preferences.
Algorithm 1 Multi-Objective Preference Optimization (MPO)
1:  Dataset: Prompt set $\mathcal{C}$, question set for reward model $\mathcal{Q}=\left\{q_{1},q_{2},...,q_{m}\right\}$
2:  Input: Diffusion model $p_{\theta}$, reward model $r$, the number of generate images per prompt $K$
@@
5:     Sample a set of images $\left\{x_{0}^{1},x_{0}^{2},...,x_{0}^{K}\right\}\sim p_{\theta}(x_{0}|c)$
6:     for $i=1$ to $K$ do
@@
8:        for $j=1$ to $m$ do
9:           Add reward $r(q_{j}|x_{i})$ to $R_{i}$
10:     for $i,j\sim\left\{1,2,...,K\right\}$ do
@@
3 Experiments
3.1 VisionReward: Preference Prediction
Dataset & Training Setting. After balanced sampling (Cf.  Sec. 2.2), we obtain 40,743 images and corresponding training data containing 97,680 judgment questions from images, leaving 6,910 images for subsequent validation and test. For videos, we obtain 28,605 videos and corresponding training data containing 89,473 judgment questions, with 3,080 videos reserved.
To fine-tune CogVLM2 [16], we set a batch size of 64, a learning rate of 1e-6, and train for 1,500 steps. For CogVLM-Video, we set a batch size of 64, a learning rate of 4e-6, and train for 1,500 steps using 32 A800 GPU hours.
To establish a comprehensive evaluation benchmark for both image and video generation, we construct MonetBench, which contains separate test sets for images and videos, each consisting of 1,000 prompts.  Tab. 3 shows content and challenge categories for MonetBench. More details are introduced in the Appendix.
Tab. 4 shows that VisionReward obtains state-of-the-state results in multiple datasets. Notably, in video evaluation, image reward models demonstrate competitive performance when the video duration is within 2 seconds (GenAI-Bench). However, when the video duration reaches 6 seconds (MonetBench), only VisionReward is capable of accurately predicting human preference, being twice (22.1% over random) as high as the best (12.5% over random) among other methods. This indicates that dynamic information in longer videos poses a challenge for RMs, while VisionReward can effectively address this issue.
@@ 3.1 VisionReward: Preference Prediction
Main Results: Accuracy and Consistency on Judgment. To evaluate the effectiveness of judgment learning, we construct a visual quality QA set to assess the visual assessment capabilities of VisionReward compared to other VLMs. We collect 1,364 test cases for images involving 14 types of questions across 4 dimensions, and additionally 1,308 cases covering 8 types of questions across 4 dimensions related to dynamic content in videos. To ensure the generality of these questions, we combine adjacent degrees in the checklists under each sub-dimension, enhancing distinctiveness and minimizing incidental subjectivity.  Tab. 5 demonstrates VisionReward’s superiority in judging visual quality.
@@
Ablation Study: Size of Train Set for Regression. We perform different scales from our training set comprising 44,000 pairs used for logistic regression. Each sampled set is then utilized for iterative regression, and the accuracy is tested on the HPDv2 test set.  Tab. 7 shows that the performance of the regression improves with the initial growth of the size of training set and then stabilizes.
Size 100 200 500 1k
Accuracy 76.5 77.6 80.3 80.6
Size 2k 4k 8k 16k
Accuracy 80.9 81.3 81.2 81.3
Table 7: Average accuracy for different regression sizes.
3.2 MPO for Text-to-Image Optimization
We maintain consistent training parameters and dataset sizes across all experiments to ensure fair comparison. For all three experiments, we used an effective batch size of 256 (with GAS set to 4 and train batch size set to 1), set $\beta$ to 5000, and a learning rate of 5e-9 (before scaling). We employ a constant warmup strategy with 100 steps. The training was conducted over 3,000 steps (approximately 1 epoch) using 64 A800 GPUs.
Experimental Results. Main results are demonstrated in  Tab. 8 and  Tab. 9. MPO with VisionReward gets leading results across multiple machine metrics and achieves significant improvements across all four dimensions of VisionReward.  Fig. 4 shows that results of MPO with VisionReward are most preferred by humans.
Methods CLIP Aes HPSv2 PickScore
Baseline 0.273 5.463 0.282 22.25
DPO with Pick-a-Pic 0.279 5.511 0.286 22.45
DPO with HPSv2 0.277 5.599 0.292 22.58
MPO (Ours) 0.279 5.612 0.289 22.61
Table 8: Evaluation results of multiple metrics on DrawBench.
Methods Composition Quality Fidelity Safety&Emotion
Baseline 0.755 0.550 0.009 -0.008
DPO with Pick-a-Pic 0.765 0.588 0.009 -0.009
DPO with HPSv2 0.874 0.630 0.010 -0.004
MPO (Ours) 0.894 0.670 0.017 -0.001
Table 9: Evaluation results analyzed by VisionReward.
Figure 4: Human evaluation of text-to-image MPO.
Ablation Study: MPO or Single RM. To evaluate the effectiveness of multi-objective versus single-objective direct preference learning, we utilize VisionReward and conduct a comparative experiment using a weighted single score as the DPO. Specifically, we sample 760k pairs under the MPO method, and sample 780k pairs using the weighted single score (threshold of 0.014). We train for 2.2k iterations with a batch size of 1,024 and a learning rate of 1e-8. The results are shown in  Tab. 10, which indicates that multi-objective can achieve better results.
Methods CLIP Aes HPSv2 PickScore
Baseline 0.273 5.463 0.282 22.25
DPO with VisionReward 0.278 5.664 0.291 22.227
MPO with VisionReward 0.278 5.719 0.291 22.505
Table 10: Evaluation results on DrawBench.
3.3 MPO for Text-to-Video Optimization.

### v2

@@ Abstract
Visual generative models have achieved remarkable progress in synthesizing photorealistic images and videos, yet aligning their outputs with human preferences across critical dimensions remains a persistent challenge. Though reinforcement learning from human feedback offers promise for preference alignment, existing reward models for visual generation face limitations, including black-box scoring without interpretability and potentially resultant unexpected biases. We present VisionReward, a general framework for learning human visual preferences in both image and video generation. Specifically, we employ a hierarchical visual assessment framework to capture fine-grained human preferences, and leverages linear weighting to enable interpretable preference learning. Furthermore, we propose a multi-dimensional consistent strategy when using VisionReward as a reward model during preference optimization for visual generation. Experiments show that VisionReward can significantly outperform existing image and video reward models on both machine metrics and human evaluation. Notably, VisionReward surpasses VideoScore by 17.2% in preference prediction accuracy, and text-to-video models with VisionReward achieve a 31.6% higher pairwise win rate compared to the same models using VideoScore. All code and datasets are provided at https://github.com/THUDM/VisionReward.
@@ 1 Introduction
Figure 1: Illustration of how VisionReward works for evaluation and optimization of visual generation. a) Evaluation: VisionReward performs comprehensive evaluation through dimension-specific binary visual QA testing, producing human-aligned, fine-grained assessment scores. b) Optimization: VisionReward enable better preference optimization, enhancing multiple key aspects.
Visual generative models, including text-to-image [9, 33, 35, 34, 2, 31] and text-to-video [17, 15, 42, 54, 4, 50] generation, have recently experienced rapid developments. Through large-scale pretraining, these models can effectively translate textual descriptions into photorealistic images or temporally coherent videos. To further align them with human preferences, reinforcement learning from human feedback (RLHF) [30]—initially introduced in large language models—has recently been adapted to visual generation tasks [49, 12].
A key bottleneck in applying RLHF to visual generation lies in developing effective visual reward models. Recent studies [49, 22, 47] have explored training reward models to predict human visual preferences, enabling automatic evaluation and preference optimization for visual generative models. For evaluation, reward models function as automated metrics that quantitatively measure the alignment between generated outputs and human preference criteria [23]. For optimization, reward models identify reliable directions for improving visual generation models. Essentially, they can provide feedback in reinforcement learning or generate preference pairs, thus reducing dependence on human annotation [3, 10, 7].
Despite recent progress in reward models (RMs) for visual generation, two primary challenges remain: First, lack of interpretability and risk of unexpected bias. Current RMs [49, 22, 47] for visual generation often suffer from limited interpretability. These models inherently involve complex trade-offs among multiple factors, yet their scoring mechanisms lack transparency regarding how such trade-offs are quantified and balanced. This opacity raises concerns about potential unexpected biases. Though multimodal LLMs like Gemini [41] and GPT-4o [1] enhance interpretability through explainable rating rationales, their general-purpose architectures usually underperform specialized black-box models in fine-grained technical assessments [6]. This raises a key dilemma: how to design new preference prediction methods that allow preference prediction to be interpretable while maintaining accuracy.
Second, lack of effective reward models for video generation. The rapid development of text-to-video generative models has intensified the demand for video reward models, yet this crucial component remains underdeveloped to date. Although image reward models can assess individual frame quality, their frame-level nature inherently neglects essential temporal dependencies in video sequences. While VideoScore [12] has pioneered direct video evaluation through learnable metrics, it still suffers from limitations such as insufficient accuracy in preference prediction and optimization in video generation.
Contribution. To address these challenges, we propose a general framework VisionReward to build accurate reward models for both image and video generation. VisionReward is trained with two steps: fine-grained visual assessment and interpretable preference learning. First, to capture human visual preferences, we identify nine major dimensions and decouple preferences into 64 fine-grained questions. Second, to ensure interpretable preference learning, we propose to use the classical linear weighting mechanism on the question outcomes. It enables intuitive visualization of each question’s impact.
To apply it as a reward model for visual generation models, we propose a multi-dimensional consistent strategy during preference optimization. The goal of this strategy is to mitigate unintended and unquantifiable biases. Specifically, a pair of visual samples is used for preference optimization (e.g., DPO [43]) only if one sample is consistently preferred over the other across all nine dimensions above.
Note that existing approaches [52, 25] have attempted to augment human annotations or expand dimensions of human preferences in visual generation. Different from them, VisionReward defines multi-dimensional human preferences with the goal of 1) disentangling distinct factors to decouple human preferences and 2) thus identifying consistently-preferred samples across all dimensions for preference optimization.
To summarize, we present VisionReward as a general framework for visual preference learning. Empirically, we show that VisionReward makes the following contributions:
  • For visual preference prediction, VisionReward achieves state-of-the-art performance across multiple benchmarks, while maintaining interpretability via hierarchical diagnostic QA and explicit linear weighting. For instance, VisionReward outperforms VideoScore [12] by 17.2% in accuracy on the human preference prediction task.
  • For visual generation, VisionReward can serve as an effective reward model for preference optimization (e.g., DPO), significantly enhancing the text-to-image and text-to-video models. For example, video generation models with VisionReward achieve a 31.6% higher pairwise win rate compared to the same models using VideoScore.
Figure 2: Overall framework of VisionReward. 1) Fine-grained Visual Assessment: fine-tune multimodal LLM to perform binary visual question-answering through hierarchical dimensions. 2) Interpretable Preference Learning: utilize visual QA outputs to predict preferences through linear weighted summation. 3) Multi-Dimensional Preference Optimization: optimization strategy across multiple dimensions.
2 VisionReward
The complete training process of VisionReward and its application methodology during preference optimization are illustrated in  Fig. 2. The training of VisionReward comprises two synergistic stages:
  • Fine-grained Visual Assessment: we develop a hierarchical framework and fine-tune the visual language model to obtain accurate visual assessment capability.
  • Interpretable Preference Learning: Leveraging the fine-tuned visual language model, we convert its results into binary entries and apply linear regression to learn the optimal weight for predicting preferences.
2.1 Fine-grained Visual Assessment
@@
Hierarchical Diagnostic Framework. When evaluating an image or video, human preferences are often a result of the interplay of multiple factors, necessitating a balance among various considerations. To deconstruct human preferences systematically, we develop a hierarchical diagnostic framework across multiple dimensions (e.g., text alignment, fidelity, safety). Each dimension is operationalized through binary questions requiring yes/no responses. There are 5 dimensions and 61 binary questions for text-to-image, while 9 dimensions and 64 binary questions for text-to-video.
Dataset Annotation. To ensure diversity in our annotated data, we sample from various data sources, preparing 48K images and 33K videos for annotation. Through specialized annotation, we obtain an image dataset containing 3 million question-answer pairs and a video dataset with 2 million pairs (Cf. Appendix  Sec. 7 and  Sec. 10 for details).
@@
Visual Instruction Tuning. Specifically, we use CogVLM2 [18] as the base model for image understanding, and CogVLM2-Video [18] as the base model for video understanding. In terms of data, we have obtained millions of annotated binary question-answering pairs. Initially, we performed a balanced sampling on each binary question by addressing the imbalance between positive and negative examples, ensuring a roughly equal number of positive and negative instances associated with each binary question. Then we use balanced instruction tuning dataset consisting of binary questions to fine-tune base VLM.
2.2 Interpretable Preference Learning
@@
Preference Learning. Based on a series of binary responses (yes or no), we use linear weighted summaries to predict human preferences. This is achieved by mapping yes and no to 1 and -1, respectively, and constructing a feature vector $\mathbf{X}=\{x^{i}\}$, where $i=1,\ldots,n$ and $n$ represents the number of binary questions. We use a set of linear weights to obtain a score that accurately reflects human preferences, expressed as follows:
@@
Through minimizing this objective function, we aim to find the optimal weights $\mathbf{w}$ to predict preferences.
Multi-Dimensional Prediction. Since weights of binary questions are get, we can define score of every question by $\text{reward}(q^{i})=w^{i}*x^{i}$, where $w^{i}$ and $x^{i}$ is the corresponding weight and feature result of question $q^{i}$. For every dimension, we can define the score of certain dimension $d_{k}$ as reward $R(d_{k})$, and $R(d_{k})=\sum\text{reward}(q^{i})$ where $q^{i}\in d_{k}$. It’s clearly that the final reward is the sum of $R(d_{k})$ from all dimensions.
After determining the weights for binary-response questions, we formulate the reward function for each question as:
$\text{reward}(q^{i})=w^{i}\cdot x^{i}$ (4)
where $w^{i}$ denotes the weight and $x^{i}$ represents the corresponding feature value for question $q^{i}$. For each evaluation dimension $d_{k}$, we define its dimensional reward $R(d_{k})$ as the summation of rewards from all constituent questions within that dimension:
$R(d_{k})=\sum_{q^{i}\in d_{k}}\text{reward}(q^{i})$ (5)
The total system reward is subsequently computed as the cumulative sum across all dimensions:
$R_{\text{total}}=\sum_{k}R(d_{k})$ (6)
Upon completion of VisionReward training, the system can execute visual diagnosis and linear weighting through a two-phase framework to predict human preferences.
2.3 Multi-Dimensional Preference Optimization
By calculating dimension-specific scores through intra-dimensional weighting, VisionReward facilitates multi-dimensional preference optimization. To empirically validate the model’s capacity, we leverage Direct Preference Optimization (DPO) for Diffusion Models in our experiments, where VisionReward generates multi-dimensional preference pairs to guide the optimization process while maintaining inter-dimensional balance.
@@ 2.3 Multi-Dimensional Preference Optimization
Challenges. Although Diffusion-DPO has designed a direct preference optimization method to fine-tune text-to-image diffusion models, it suffers from being biased and over-optimized on certain dimensions such as aesthetic but gets somewhat worse, for example, less safe and more trouble about human body. We reproduce the procedure of Diffusion-DPO and test the fine-tuned model on 10,000 sampled prompts from training data, using VisionReward to score images.  Fig. 3 illustrates the changes across various dimensions after applying preference learning using the human preference dataset in which factors are coupled with different trade-offs.
MPO: Insight and Solution. We propose a multi-dimensional preference strategy (see Algorithm 1) to ensure comprehensive enhancement across all dimensions. Following Sec. 2.2, we obtain dimensional reward $R(d_{k})$ and total reward $R_{\text{total}}$ for each generated image. For two images $x^{i}$ and $x^{j}$, we define $R^{i}$ as dominating $R^{j}$ if $R(d_{k}^{i})\geq R(d_{k}^{j})$ holds for every dimension $d_{k}$. The key differences between the MPO strategy and standard DPO are:
  • Standard DPO: During DPO optimization, we generate $m$ images per prompt and directly select the pair with the maximal difference in total reward $R_{\text{total}}$.
  • MPO-enhanced DPO: MPO strategy introduces an additional constraint: we only select pairs $x_{0}^{i}$ and $x_{0}^{j}$ if $R^{i}$ dominates $R^{j}$, then proceed with standard DPO.
We analyze the effects of MPO in Sec. 3.3. The MPO strategy can also be applied to other algorithms, which we leave for future exploration.
Algorithm 1 Multi-Dimensional Preference Optimization (MPO)
1:  Dataset: Prompt set $\mathcal{C}$, question set for reward model $\mathcal{Q}=\left\{q_{1},q_{2},...,q_{n}\right\}$
2:  Input: Diffusion model $p_{\theta}$, reward model $r$, the number of generate images per prompt $m$
@@
5:     Sample a set of images $\left\{x_{0}^{1},x_{0}^{2},...,x_{0}^{m}\right\}\sim p_{\theta}(x_{0}|c)$
6:     for $i=1$ to $m$ do
@@
8:        for $j=1$ to $n$ do
9:           Add reward $r(q_{j}|x_{0}^{i})$ to $R_{i}$
10:     for $i,j\sim\left\{1,2,...,m\right\}$ do
@@
3 Experiments
3.1 VisionReward for Text-to-Vision Evaluation
Dataset & Training Setting. After balanced sampling, we obtain 40,743 images and corresponding training data containing 97,680 judgment questions from images, leaving 6,910 images for subsequent validation and test. For videos, we obtain 28,605 videos and corresponding training data containing 89,473 judgment questions, with 3,080 videos reserved.
To fine-tune CogVLM2 [18], we set a batch size of 64, a learning rate of 1e-6, and train for 1,500 steps. For CogVLM-Video, we set a batch size of 64, a learning rate of 4e-6, and train for 1,500 steps.
To establish a comprehensive evaluation benchmark for both image and video generation, we construct MonetBench, which contains separate test sets for images and videos, each consisting of 1,000 prompts.  Tab. 1 shows content and challenge categories for MonetBench. More details are introduced in Appendix  Sec. 11.
Tab. 2 shows that VisionReward obtains state-of-the-state results in multiple datasets. Notably, in video evaluation, image reward models demonstrate competitive performance when the video duration is within 2 seconds (GenAI-Bench). However, when the video duration reaches 6 seconds (MonetBench), only VisionReward is capable of accurately predicting human preference, being twice (22.1% over random) as high as the best (12.5% over random) among other methods. This indicates that dynamic information in longer videos poses a challenge for RMs, while VisionReward can effectively address this issue after fine-grained visual learning.
Main Results: Accuracy and Consistency on Judgment. To evaluate the effectiveness of judgment learning, we construct a visual quality QA set to assess the visual assessment capabilities of VisionReward compared to other VLMs. We collect 1,364 test cases for images involving 14 types of questions across 4 dimensions, and additionally 1,308 cases covering 8 types of questions across 4 dimensions related to dynamic content in videos. To ensure the generality of these questions, we combine adjacent degrees in the checklists under each sub-dimension, enhancing distinctiveness and minimizing incidental subjectivity.  Tab. 3 demonstrates VisionReward’s superiority in judging visual quality over existing generalized multimodal LLMs, which remains a challenge for generalized multimodal LLMs.
@@ 3.1 VisionReward for Text-to-Vision Evaluation
Ablation Study: Impact of Training Set Size. We conduct experiments with varying sizes of the training set to investigate its influence on regression performance. As demonstrated in  Tab. 5, the accuracy improves monotonically as the training set size increases up to 4,000 samples, beyond which the accuracy stabilizes.
Ablation Study: Scalability of Question Scale.  Tab. 6 shows that accuracy of preference prediction exhibits significant improvement as the question scale increases. Scalability validates the effects of decomposing preferences via multi-scale questioning. We do more analysis such as question pruning in Appendix  Sec. 8.
Size 200 500 1k 2k 4k 8k
Acc. 77.6 80.3 80.6 80.9 81.3 81.3
Table 5: Accuracy on HPDv2-test for different sizes of train set.
#question 2 4 8 16 32 all
Acc. 59.3 68.6 72.5 75.8 78.9 81.7
Table 6: Accuracy on HPDv2-test for different question scales.
3.2 VisionReward for Text-to-Vision Optimization
Figure 4: Human evaluation results of text-to-image DPO.
Figure 5: Human evaluation results of text-to-video DPO.
To evaluate the efficacy of VisionReward in preference optimization for visual generative systems, we conducted a series of comparative experiments against current state-of-the-art reward models and established preference datasets. The empirical results presented in  Fig. 4 and  Fig. 5 demonstrate VisionReward’s superior performance in both text-to-image and text-to-video optimization, achieving statistically significant improvements in human preference metrics over competing approaches.
This section focuses on preference optimization for text-to-video using different reward models. Details of text-to-image are provided in Appendix  Sec. 9.2 due to page limits.